OpenStudy (anonymous):

how much would $500 dollars invested at 7% interest compounded annually be worth after 4 years? Round your answer to the nearest cent.

OpenStudy (anonymous):

I got $655.40

OpenStudy (anonymous):

is that rite

OpenStudy (anonymous):

For 4 Year's.

OpenStudy (anonymous):

yeah 4 years

OpenStudy (anonymous):

I got 700.00

OpenStudy (anonymous):

I did 500 times 7% times 4.

OpenStudy (anonymous):

well you'd use the compound interest formula of course

OpenStudy (anonymous):

im just checking if my answer is correct

OpenStudy (anonymous):

hello?

OpenStudy (anonymous):

yeah & I got $655.40

OpenStudy (anonymous):

is that correct

OpenStudy (anonymous):

I / pt = r

OpenStudy (anonymous):

I = 0.07

OpenStudy (anonymous):

p = 500

OpenStudy (anonymous):

t = 4

OpenStudy (anonymous):

p * t or 500 * 4 =2000

OpenStudy (anonymous):

yea I did all this but I just want to make sure if my answer is right

OpenStudy (anonymous):

if you followed the procedure i guest it would be right

OpenStudy (anonymous):

i used the compound interest formula

OpenStudy (anonymous):

anyways what abt this? You deposit $425 into the account, where it earns 6.2% annual interest, compounded monthly, and leave it there for 18 months.

OpenStudy (anonymous):

same type of problem but im a little confused on what u do w/ the 18 months

OpenStudy (anonymous):

you turn 18 months to years 1.6 years

OpenStudy (anonymous):

oh ok. wouldn't it be 1.5years though

OpenStudy (anonymous):

cuz it's 1 1/2 mos.

OpenStudy (anonymous):

sorry 1 1/2 yrs

OpenStudy (anonymous):

theres 12 months in a year so 1 year and then 18 m - 12 m = 6 m so 1 y and 6 m